OpenStudy (hawkfalcon):

simplify sinθ*sec^2(90-θ)

OpenStudy (hawkfalcon):

Simplify

OpenStudy (hawkfalcon):

\[\sinθ*\sec^2(90-θ)\]

OpenStudy (dumbcow):

using the sum/difference angle identities you can see that \[\cos(90-\theta) = \cos(90)\cos(\theta) + \sin(90)\sin(\theta) = \sin(\theta)\] thus \[\sec^{2}(90-\theta) = \csc(\theta)\]

OpenStudy (dumbcow):

sorry that should be csc^2

OpenStudy (noelgreco):

sec(90-theta) is the secant of the COmplement of theta or the COsecant theta.

OpenStudy (hawkfalcon):

Oh okay:3 that makes sense. Thank you!:D